# ChauBoxplot and AdaptiveBoxplot: two R packages for boxplot-based outlier detection

### v1

@@ 5 Conclusion and Practical Guidance
In this paper, we reviewed recent advances in boxplot-based outlier detection, focusing on the ChauBoxplot and AdaptiveBoxplot R packages. Our simulation results (see Figure 1) demonstrate that the classic Tukey boxplot increasingly misclassifies regular data points as outliers as the sample size increases. In contrast, the Chauvenet-type boxplot consistently identifies the true outliers in large datasets (e.g., n=50,000n=50,\!000), while minimizing false positives. The FDR-controlled boxplot is also sensitive to outliers, though it may occasionally miss some extreme cases. Meanwhile, the FWER-controlled method is highly conservative and can sometimes overlook genuine outliers, particularly in large samples.
Based on extensive simulation studies (including many results not shown here due to space limitations), we recommend any boxplot method for small datasets, with the FWER-based approach being particularly suitable due to its conservativeness. For moderate-sized datasets (typically 100 to 1,000 observations), the Chauvenet-type and FDR-based boxplots provide the best balance of sensitivity and specificity. In large datasets, the fixed-threshold rules should be avoided; the Chauvenet-type boxplot remains robust across all sample sizes, while FDR-based methods—with a relaxed significance level—offer greater sensitivity when multiple outliers are expected. For all data sizes, robust estimation and careful consideration of the underlying distribution are essential, especially with skewed or non-normal data. Supported by well-documented R packages, these modern approaches enable more accurate, interpretable, and reproducible outlier detection, thereby providing a solid foundation for future advances in statistical graphics and robust data analysis.

### v2

@@ 6 Conclusion and practical guidance
This paper reviewed recent advances in boxplot-based outlier detection, focusing on the ChauBoxplot and AdaptiveBoxplot R packages. Results from simulations and real-world data demonstrate that Tukey’s classic boxplot, with its fixed-fence rule, increasingly misclassifies central observations as outliers as the sample size grows, while the sample-size-adjusted and error-controlled approaches, namely the Chauvenet-type, Holm-type, and BH-type boxplots, offer substantially improved reliability. The Chauvenet-type boxplot provides explicit control over the expected number of flagged points, the Holm-type approach controls the probability of at least one false outlier, and the BH-type method adapts flexibly while bounding the false discovery rate. Table 2 provides a concise summary of the error control and practical characteristics of these methods.
Method Error Control Flagging Behavior
Tukey’s boxplot None (fixed fence) Flags a fixed proportion as outliers under normality; the rate increases rapidly with the sample size.
Chauvenet-type boxplot Expected number (PFER) Flags on average 0.5 outlier per dataset, regardless of the sample size.
Holm-type boxplot Family-wise error rate (FWER) Controls the probability of at least one false outlier (typically $\leq$5%).
BH-type boxplot False discovery rate (FDR) Controls the expected proportion of false outliers among those flagged (typically $\leq$5%).
Table 2: Summary of boxplot outlier detection methods and their error interpretation. PFER: per-family error rate (expected false outlier count); FWER: family-wise error rate (probability of at least one false outlier); FDR: false discovery rate (expected proportion of false outliers among those flagged).
Our simulation studies and practical analyses, as illustrated in Section 5 with $n=31$, reveal that the choice of boxplot method can influence which outliers are detected, even in relatively small samples. The classical Tukey boxplot and the Chauvenet-type boxplot sometimes yield overlapping but not identical sets of flagged values, especially when the data deviate from normality or exhibit moderate skewness. For small datasets, both traditional and adaptive boxplot methods are generally reasonable; however, users should consider the distributional characteristics of their data and interpret results with caution. The Holm-type approach remains particularly appropriate in applications demanding strict control of false positives. For moderate sample sizes, typically 100 to 1000 observations, the Chauvenet-type and BH-type boxplots offer the best balance of sensitivity and specificity. In large datasets, the fixed-threshold rules should be avoided, as their rate of false positives escalates. The Chauvenet-type boxplot remains robust across all sample sizes, while the BH-type method with a relaxed significance level provides greater sensitivity when multiple outliers are expected.
It is important to recognize, however, that most boxplot-based techniques, including both ChauBoxplot and AdaptiveBoxplot, are developed under the assumption of normality. As illustrated by the real data example in Section 5, real-world datasets such as the provincial Project 985 and Project 211 admission rates can display marked skewness, heavy tails, or other departures from normality due to social, economic, or policy-driven factors. In these cases, even sample-size-adjusted and error-controlled boxplots may produce misleading results. For example, skewed data can result in fences that are either overly strict or overly lenient, which may cause atypical observations to be missed or typical values to be incorrectly labeled as outliers. Practitioners are encouraged to assess distributional characteristics before applying these methods and to consider robust estimation or data transformations as appropriate. Future research directions include developing adaptive boxplot procedures that remain valid under a wider array of data distributions, potentially through transformation-invariant fences or alternative criteria tailored to skewed or heterogeneous datasets.
In summary, our simulation studies and real-world analysis demonstrate that ChauBoxplot and AdaptiveBoxplot provide reproducible, statistically principled tools for modern boxplot-based outlier detection across a wide range of applications. By embedding adaptive error control into familiar graphical displays, these packages empower practitioners to achieve defensible and interpretable results in both exploratory and confirmatory analyses. We recommend these approaches particularly for moderate to large datasets, and emphasize the importance of assessing distributional characteristics to ensure robust and reliable inference.
Acknowledgement(s)
The authors sincerely thank the Editor and the two reviewers for their thoughtful and constructive comments, which have greatly improved this article.
Funding
Tiejun Tong’s research was supported in part by the General Research Fund of Hong Kong (HKBU12300123) and the Initiation Grant for Faculty Niche Research Areas of Hong Kong Baptist University (RC- FNRA-IG/23-24/SCI/03). Hongmei Lin’s research was supported in part by the National Natural Science Foundation of China (12171310). Riquan Zhang’s research was supported in part by the National Natural Science Foundation of China (12371272, 12531013).